OpenStudy (miandra27):

What steps do you take to factor the following expressions? x^2 - 4 x^3 - 27

TheSmartOne (thesmartone):

The same does not apply to the second question. Use this formula for the second question: \(\Large a^3 + b^3 = (a+b)(a^2-ab+b^2)\) \(\Large a^3 - b^3 = (a-b)(a^2+ab+b^2)\)

OpenStudy (sshayer):

\[a^2-b^2=(a+b)(a-b)\]